Find the radius and the centre of the circle x^2+y^2+z^2-x+z-2=0, x+2y-z=4
1
Expert's answer
2020-06-15T19:36:54-0400

x²+y²+z²+-x+z-2=0

x+2y-z==4


add the equations x²+y²+z²+-2y-6=0

or x²+y²+2y+1+z²-7=0

x²+(y+1)²+z²=7


hence

circle center (0, -1.0)

circle radius √7